Division of Spill Prevention and Response

Breadcrumbs

Chronology

SITUATION UPDATE – December 5, 2009

ACTIVITIES

  • Approximately 40 cubic yards of contaminated material manually removed bringing the total volume to approximately 160 cubic yards. Material will be stored at the East Dock ice pit until it is melted and the volume of oil measured.
  • Analysis of the initial X-ray survey images indicates extensive ice plugs within the 18-inch line; one ice plug measured approximately 1,500 feet in length.
  • The 40-foot diameter exclusion zone may be retracted, allowing access to the heavily contaminated product closest to the location of the pipeline leak, with additional safety considerations.
  • Planning underway to employ hot steam (steamer) tactics with vacuum truck recovery for site cleanup. The steamer tactic applies localized heat to the snow and oiled material and the resultant liquid product is recovered by the vacuum truck. This tactic generates less waste and is more effective than water flushing. Steamer tactics will commence once all necessary equipment is moved to the staging area.
  • A 3-foot ice berm was constructed around the spill site to contain oily water generated by the steamer tactic.

FUTURE PLANS AND RECOMMENDATIONS

  • Steamer operations to continue until all gross contamination has been collected
  • Water flush and vacuum recovery tactics may also be applied to the site if the steamer tactic is not totally effective. Warm-water flushing melts the contaminated snow and washes the tundra. The contaminated water is then collected in vacuum trucks.

return to top

SITUATION UPDATE � December 6, 2009

ACTIVITIES

  • Crews manually removed an additional estimated 40 cubic yards of contaminated material bringing the total volume of contaminated material removed from the site to approximately 200 cubic yards. Material will be stored at the East Dock ice pit until it is melted and the volume of collected oil measured. 
  • All necessary steamer equipment has been moved to the staging area. BPXA is developing a job hazard assessment for Unified Command (UC) approval to ensure worker safety.   Steaming will commence once safety and operational issues are addressed.
  • The initial ice berm around the spill site was completed.  Additional water was added to the berm overnight to improve integrity.  However, an increase in ambient temperature at Prudhoe Bay has impacted the ability to make additional ice.  The use of the steamer tactic will not be affected if the ice berm cannot be improved as contaminated liquid will be collected after each application of steam, rather than allowing liquid to build up within the containment area.
  • Work on the ice road continues between the Lisburne Production Center and the South Division operational area.  The ice ramp and turn around area at the spill site has been completed.

FUTURE PLANS AND RECOMMENDATIONS

  • Once initiated, steamer operations will continue until all gross contamination has been collected. 
  • If the steamer tactic is unable to adequately collect contamination, water flush and vacuum recovery tactics may be applied to the site.  This tactic will require further reinforcement of the ice berm. 
return to top

SITUATION UPDATE – December 7, 2009

ACTIVITIES

  • The steamer tactic began yesterday at approximately 4:00 PM.  Work crews collected approximately 3,300 gallons of contaminated material using this cleanup method during the night shift that ended at 6:00 AM today; this includes the approximately 290 gallons of water resulting from the steam used as part of the cleanup tactic.
  • Warm-water flushing in a natural low spot has also been initiated within the heavily contaminated area. Work crews apply controlled amounts of water to the contaminated area, and vacuum equipment is used to collect the product immediately, allowing different cleanup operations to continue at other locations within the ice berm.
  • Work crews manually removed approximately 50 cubic yards of contaminated material during the 24-hour operational period that ended at 6:00 AM today. This material will be stored at the East Dock ice pit until it is melted, combined with collected liquid product, and the volume of collected oil measured.
  • Approximately 0.4 miles of the ice-road has been completed. The remainder of the ice-road has been roughed in and is awaiting colder weather for completion. BPXA is exploring other options to get equipment on-site; the Unified Command has approved moving limited equipment along the packed snow base of the ice-road. Oiled material will be collected mechanically once equipment is on site.

FUTURE PLANS AND RECOMMENDATIONS

  • Steamer and water-flush cleanup tactics will continue until determined to be no longer necessary.
  • Mechanical removal of oiled material will commence when equipment is on site. Once mechanical operations are underway, the manual removal of snow is expected to cease.
return to top

SITUATION UPDATE – December 8, 2009

ACTIVITIES

  • Work crews recovered approximately 3,738 gallons of contaminated material using the steamer-head during the 24-hour operational period that ended at 6:00 AM today. Steaming operations have essentially been stopped because of safety concerns associated with overlapping cleanup tactics.

  • A bobcat rubber-track loader was moved within the containment area and is mechanically collecting contaminated material. The material is being collected in totes and passed to clean equipment outside the contamination zone for ultimate transfer to the East Dock for storage. Approximately 64 cubic yards of contaminated material was mechanically removed from the site during the night shift that ended at 6:00 AM today.

  • Warm-water flushing has ceased since the mechanical snow removal operations commenced. Plans are primarily to use mechanical snow removal tactics, but flushing equipment will be kept on site as a backup.

  • Work crews manually removed approximately 65 cubic yards of contaminated material during the 24-hour operational period that ended at 6:00 AM today. The total quantity of manually removed material is approximately 325 cubic yards. The material has been transported and temporarily stored at the East Dock.

  • Work on the ice-road continues; approximately 0.9 miles of the ice-road is complete. Contaminated snow is transported across the tundra using rolligon equipment. The rolligon will be replaced with flatbed trucks once the ice-road is complete. A rolligon is a specialized, low-impact, off-road vehicle.

FUTURE PLANS AND RECOMMENDATIONS

  • Manual and mechanical snow removal will continue until it is determined to be no longer necessary for cleanup operations. Direct access to the spill site will be possible once the ice-road is completed.
return to top

SITUATION UPDATE – December 9, 2009

ACTIVITIES

  • A bobcat rubber-track loader is mechanically collecting contaminated material from within the spilled area. The material is collected in totes, passed to clean equipment outside the contamination zone, and ultimately transferred to the East Dock for storage.   Mechanical recovery continues to be the primary clean-up method.
  • Work crews mechanically removed approximately 226 cubic yards of contaminated from the site during the 24-hour operational period ending at 6:00 AM today.
  • Steaming operations have ceased; equipment for this operation was demobilized, but could be recalled within 24 hours. The volume of contaminated material recovered through this tactic is 6,174 gallons.
  • Work on the ice-road continues; approximately 1.1 miles (or approximately 90%) of the ice-road is complete.  BPXA continues the moving of contaminated snow across the tundra using rolligon equipment.  A rolligon is a specialized, low-impact, off-road vehicle.  The rolligon will be replaced with flatbed trucks once the ice-road is complete. 

FUTURE PLANS AND RECOMMENDATIONS

  • Mechanical and manual snow removal will continue until determined to be no longer necessary.   The colder weather should help with the completion of the ice-road, allowing direct access to the spill site.
return to top

SITUATION UPDATE – December 10, 2009

ACTIVITIES

  • The Unified Command Safety Officer has determined that mechanical equipment is now allowed within the 10-foot exclusion zone around the pipeline.
  • Mechanical and manual cleanup of the contaminated area continued through the night.  Work crews removed approximately 173 cubic yards of contaminated material in the 24-hour operational period that ended at 6:00 AM today.
  • Water flushing and vacuum removal of contaminated fluids resumed at 2:30 AM today. Cleanup workers are using warm water on heavily contaminated areas, flushing and recovering contaminated liquids in a topographic low point. An ice and snow containment berm is maintained around the spill area to prevent the spread of contaminants.  As of 6:00 AM today, the vacuum trucks have removed an estimated 156 barrels of liquid.
  • The Unified Command approved site access, after assessing and mitigating the safety hazards, for agency representatives to obtain photos of the 18-inch pipeline leak site.

FUTURE PLANS AND RECOMMENDATIONS

  • Mechanical and manual snow removal will continue until other tactics are approved by the Unified Command.
  • Options for repairing the pipeline are under development.
return to top

SITUATION UPDATE – December 11, 2009

ACTIVITIES

  • Cleanup workers halted water flushing operations around 11:30 PM on December 10th when a softening of the ice berm developed.  Cleanup workers inspected the ice berm, and the local and state agency representatives confirmed that the ice berm was not breeched. Cleanup workers reinforced the ice berm with more snow and water before resuming operations.
  • Work crews using mechanical and manual cleanup methods removed approximately 81 cubic yards of contaminated snow and 270 barrels of contaminated liquid as of 6:00 AM today. 

FUTURE PLANS AND RECOMMENDATIONS

  • Mechanical and manual operations will continue to remove spilled product and contaminated material.
  • Options for repairing the pipeline are under development.
return to top

SITUATION UPDATE – December 12, 2009

ACTIVITIES

  • As of 6:00 AM, December 12, cleanup workers had removed the last of the contaminated snow and vacuumed the remaining liquids. 
  • Oily liquids are being transported to Flow Station 2 for oil-water separation.
  • The ice road has been completed between the Lisburne Production Center and the spill site. 
  • Response crews have completed Phase 2 of the cleanup effort.  The cleanup is entering Phase 3, which addresses embedded oil and salt water removal.
  • Today, work crews will begin using a rotating mechanical trimmer mounted on a Bobcat to remove the topmost layer of any contaminated tundra.  Due to snow cover, and more falling snow on December 12, workers will continue to use shovels and brooms to remove snow to uncover the frozen tundra layer.

FUTURE PLANS AND RECOMMENDATIONS

  • Continue Phase 3 operations of the cleanup to remove spilled product and contaminated material.
  • Options for repairing the pipeline are under development.
return to top

SITUATION UPDATE – December 13, 2009

  • No update for December 13
return to top

SITUATION UPDATE – December 14, 2009

ACTIVITIES

  • Response crews completed Phase 2 of the cleanup effort , which included manually removed the remaining contaminated snow in the zone where no water-washing occurred.  Mechanical and manual operations generated an estimated 1,084 cubic yards of contaminated snow, which will be mechanically melted.  Snow melting operations are scheduled to begin today, with an expected completion date of December 16.
  • Cleanup workers completed the water-flushing and vacuuming of released material, recovering approximately 1,387 barrels (58,254 gallons) of oily-liquids.  The volume reported includes fluids released from the pipeline, melted snow, and water used during the flushing operation.
  • As part of Phase 3 of the cleanup plan, heavy equipment operators are using a rotating mechanical “trimmer” to remove contaminated ice and embedded oil from the tundra.  As of 6:00AM on December 14, an estimated 110 cubic yards of “trimming” solids have been removed and taken to a temporary storage cell.
  • On December 14, Bell Surveys Inc. began a site survey of the spill area that will include topographical features, locations of various spill response tactics, travel and access routes, and a site grid.  The survey will assist in determining the effectiveness of treatment and preparing plans for future monitoring of the site.
  • On December 12, crews completed the ice road between the Lisburne Production Center and the spill site, allowing alternative highway access to the pipeline and spill response area.  Work continues on the ice road to extend it to LO3 of the pipeline.

FUTURE PLANS AND RECOMMENDATIONS

  • Continue Phase 3 cleanup operations to remove remaining contaminated material, including embedded oil and saltwater from the tundra.
  • Pipeline repair plans remain under development and will be discussed with ADEC and EPA later this week.
return to top

SITUATION UPDATE – December 17, 2009

ACTIVITIES

  • Response crews have completed the bulk of the cleanup effort; officials have provided a 95% completion estimate.  The final night shift occurred December 16-17.
  • The Unified Command has stood down, marking the transition from an emergency response phase to a project phase.  ADEC continues to monitor the cleanup activities and waste management aspects of the cleanup project and investigate the cause and other aspects of the release.
  • Severe cold weather has halted cleanup efforts at times.  Workers continue to use jackhammers to remove frozen contaminated materials from beneath the pipeline.  Snow melting operations slowed down due to mechanical problems, but will continue until all contaminated snow has been liquefied and the oil-water mixture stored in a tank to await final volume recordings for oil content.
  • Oasis Environmental consultants continue to collect screening samples of the spill site.  The sampling has located a few spots that will require more mechanical operations to remove the remaining contamination.
  • ADEC has approved the confirmation sampling plan, which outlines how soil testing will confirm contaminants are properly removed in preparation for tundra rehabilitation next spring.
  • Work is almost complete on the ice road extension from the spill site to LO3; once finished, the ice road will provide alternative access to the pipeline from LPC to LO3.  High-powered “Scan-track” X-ray operations continue along the pipeline.

FUTURE PLANS AND RECOMMENDATIONS

  • Mechanical and manual cleanup operations to remove the contaminated material from hot spots will continue.  Snow melting operations will also continue.
  • Engineers are completing their review and risk assessment for the pipeline repair plan.  Repair parts are scheduled to arrive late next week.
  • With the transition to the cleanup project phase, ADEC staff are beginning to demobilize.  ADEC will maintain an appropriate field presence to monitor final cleanup, waste management, investigation, and pipeline repair operations.
return to top